### v1

# Augmenting LLMs to Repair Obsolete Test Cases with Static Collector and Neural Reranker

Abstract
During software evolution, it is advocated that test code should co-evolve with production code. In real development scenarios, test updating may lag behind production code changing, which may cause the project to fail to compile or bring other troubles. Existing techniques based on pre-trained language models can be adopted to repair obsolete tests caused by such unsynchronized code changes, especially syntactic-related ones. However, the lack of target-oriented contextual information affects repair accuracy on large-scale projects. Starting from an obsoleted test, the key challenging task is precisely identifying and constructing Test-Repair-Oriented Contexts (TROCtx) from the whole repository within a limited token size.
In this paper, we propose Synbciatr (Syntactic-Breaking-Change-Induced Automated Test Repair), a novel approach to automatically repair obsolete test cases via precise and concise TROCtx construction. Inspired by developers’ programming practices of the task, we design three types of TROCtx: class contexts, usage contexts, and environment contexts. For every type of TROCtx, Synbciatr automatically collects the changed-token-related code information through static analysis techniques. Then it generates reranking queries to identify the most relevant TROCtxs, which will be taken as the repair-required key context and be input to Large Language Model for the final test repair.
To evaluate the effectiveness of Synbciatr, we construct a benchmark dataset that contains diverse syntactic breaking changes. The experimental results show that Synbciatr outperforms baseline approaches both on textual- and intent-matching metrics. With the augmentation of TROCtx constructed by Synbciatr, hallucinations are reduced by 57.1%.
Index Terms:
Software Evolution, Test Repair, LLM, Static AnalysisI Introduction
Software evolution is a fundamental and significant aspect of software development [1]. For large-scale software such as Kafka [2], the project frequently evolves, where a new version is released in one to two weeks on average and new commits are submitted nearly every day. As production code usually changes during software evolution, it is crucial to maintain and co-evolve associated test codes to ensure that they remain effective in validating the software’s functionality [3], [4]. Specifically, for software that is invoked as libraries, it is important to co-evolve the test code to follow the changes of production code, which can help developers quickly notice the backward incompatible changes that may affect its clients [5].
To automatically co-evolve production code and test code, previous studies analyze and mine the software code to extract production-test co-evolution rules and patterns [6], [7]. However, as real-world code changes come in a great many forms, they are hard to summarize into a small number of general patterns. In recent years, with the rapid advancement of machine learning and large language models (LLM) [8], [9], many studies have utilized learning-based techniques to assist the production and test code co-evolution, including obsolete test cases identification [10], production-test co-evolution pairs extraction [11], etc., which yielded favorable results. To further reduce the developer’s burden, researchers also try to repair the obsolete test cases automatically. For example, Hu et al. [12] identify and update obsolete test methods by fine-tuning a pre-trained model initialized from CodeT5 [13], which is currently the SOTA approach for repairing obsolete test cases.
Though the learning-based techniques resulted in positive outcomes on many tasks, they still faced with difficulties when operating code in complex repositories. For example, API signature changing is the most common and straightforward code-changing type in software evolution. However, focusing on this type of production change, the SOTA work Ceprot [12] failed to correctly repair about three-quarters of test cases on real-world projects (refer to Section V-A), which means there still have a gap before it can be applied to co-evolve tests in practice.
To fill in the gap, we mainly target precisely repairing these signature-related code changes based on the power of LLMs. Here, the signature-based code changes are also called Syntactic Breaking Changes (SynBC) as they may lead to compilation errors if their associated tests are not co-evolved. According to our investigation, accurately repairing synBC-related obsolete test cases based on LLMs still faces the following key challenges.
C1:
The repair-oriented code contexts are unclear. When developers manually repair a test case, not only the changed signature of the focal method but also many other changed-token-related code contexts are required. When fixing tests with learning-based approaches, the repair-oriented code context must be explicitly extracted. As existing works repair test methods solely with the original and updated focal methods as input, it is still unclear which type of contexts are essential for accurate repair and how to extract them.
C2:
The token size of code contexts is limited. Compared to small-scale models, LLMs have demonstrated their extraordinary capabilities. However, even though they are designed with increasingly larger context windows, it is impractical to simply include the entire repository as input. And extra irrelevant information may bring negative effects as well. That means the extracted code context is not the more the better. As the number of context tokens processed in a single interaction is restricted, for all the code contexts that have a relation with the changed signature, it is required to sort and pick out the most relevant ones as part of the final context.
To this end, we propose Synbciatr (Syntactic-Breaking-Changes-Induced Automated Test Repair), a novel approach to repairing obsolete test cases caused by SynBC at the method level. To address challenge C1, Synbciatr designs three types of contexts, including Class Contexts (ClassCtx), Usage Contexts (UsgCtx), and Environment Contexts (EnvCtx) as Test-Repair-Oriented Contexts (TROCtx). These contexts focus on different aspects of changed code, but all aim at providing adequate contextual information for test-repairing together. For challenge C2, Synbciatr identifies and constructs qualified TROCtx from the repository in two stages. ❶ First, Synbciatr collects all types of TROCtx by static code analysis, specifically, by using Language Server. ❷ Then, inspired by the idea of retrieval-augmented generation (RAG) [14], Synbciatr generates reranking queries to identify the most relevant TROCtxs according to the repair requirement extracted from the original test case. That process is based on Neural Rerankers [15], a set of neural models for reranking in information retrieval. After constructing TROCtxs, Synbciatr aggregates test-repair-required information to a full LLM prompt and finally generates the repaired test cases.
To evaluate the effectiveness of Synbciatr, we construct a benchmark dataset based on existing work which consists of 136 samples with diverse SynBCs. The evaluation is based on both the textual match and intent match metrics. In terms of the textual match, Synbciatr achieves the best performance against baselines specifically on CodeBLEU (83.3), DiffBLEU (46.7), and Accuracy (32.4). In terms of the intent match, we conduct a human evaluation on verifying whether test cases are correctly repaired without changing their original intent, in which Synbciatr achieve improvements of 248.6% and 9.8% when compared to Ceprot and naiveLLM respectively. Moreover, Synbciatr is capable of reducing 57.1% hallucinations.
We make the following contributions in this paper:
We design three types of TROCtx to provide adequate contextual information for test repair caused by SynBC.
We propose Synbciatr to construct TROCtx by combining static collector and neural rerank, and use it to repair test based on LLM.
Experimental results on the benchmark dataset demonstrate that Synbciatr can repair obsolete test cases caused by SynBC more effectively compared to both Ceprot and Naivellm.
The data and code are both publicly available at: https://anonymous.4open.science/r/SynBCIATR-3DD2.
II Background and Motivation
II-A Task Definition
Referring to previous studies [16], [17], [5], Syntactic Breaking issues (SynB) represent signature-based errors in API evolution, such as ClassNotFoundException and NoSuchMethodError. In this paper, we name method changes that may cause SynB as Syntactic Breaking Changes (SynBC), in which the signature of the method changes during evolution.
We define the signature of a method (Method Signature) as a tuple , where:
is the name of the method;
is a list of parameter types;
is the type of the return value;
is a set of modifiers;
is a set of exception types that can be thrown.
Given a method that changes from to , we define Syntactic Breaking Changes (SynBC) as:
where and denote the method signatures before and after the change respectively. To figure out whether SynBCs are common in production-test co-evolution, we conducted an empirical study to count the SynBCs exist in the dataset released by Ceprot [12]. Each sample in the dataset contains a change of focal method on a real-world commit from GitHub. It reveals that signature-based focal changes occur in over 40% samples of the dataset, in which parameter and return type-related changes account for a large proportion.
Thus, we further categorize SynBC into the following three types according to the changed elements in a method signature. Note that, a SynBC whose parameter and return types in the method signature both changes belongs to ParamSynBC and RetSynBC at the same time.
Parameter-related Syntactic Change (ParamSynBC). For a SynBC from to , if the parameter type list in the method signature changes, it is a ParamSynBC.
Return-related Syntactic Change (RetSynBC). For a SynBC from to , if the type of the return value in the method signature changes, it is a RetSynBC.
Normal Syntactic Changes (NormSynBC). For a SynBC from to , if it does not belong to ParamSynBC or RetSynBC, it is a NormSynBC.
As repairing tests caused by SynBC is a primary work for effective production-test co-evolution, we make efforts to repair obsolete test cases under this type. Here, for a SynBC that happens between the original focal method and the updated focal method , denotes the repository code that method belongs to, and denotes the obsolete test case associated with . Based on this knowledge, we hope to get the updated test case as the final output. The whole task consists of two main steps. First, we extract the repair-oriented contexts for test from repository according to the SynBC from to . That process can be denoted as:
where represents the constructed contexts. Then, with the contexts , repairing obsolete tests can be expressed as:
where denotes the final repaired test case.
II-B Language Server and Neural Reranker
Language Server consists of programming language specific tools like static analyzers and compilers. In modern Integrated Development Environments (IDEs), Language Server provides language-specific features like ‘autocomplete’, ‘goto definition’, ‘find usages’, and others [18]. To help IDEs communicate with Language Server [19], Microsoft has created a standard JSON-RPC-based protocol, called Language Server Protocol (LSP), based on which multiple IDEs can communicate with the same Language Server to access intelligence features.
Neural Reranker is a type of machine learning models [20] used to reorder a given set of documents based on their relevance to the query, which is the initial request for information expressed as keywords or complex expressions [15]. It is widely used in the context of information retrieval, natural language processing, and recommendation systems. Instead of ranking based on simple heuristics like the frequency of query terms appearing in the query, Neural Rerankers are trained to take into account more complex features, like the semantic similarity between the query and the document.
II-C Motivating Example
The example in Fig. 1 is used to demonstrate our motivation. It was collected from a real commit in project Alluxio [21]. As shown in the given commit, there is a ParamSynBC for the focal method (named mount), where the third parameter changes from MountOptions to MountPOptions. If the associated original test case (also named mount) does not co-evolve with the change of focal method, the test will fail to be built with compilation error (cannot resolve type).
To automatically repair the obsolete test method, we first apply existing learning-based techniques directly. However, both Ceprot and GPT-4 fail due to using an undefined method (MountPOptions.defaults()) to construct the third parameter, where the correct method invocation should be MountPOptions.getDefaultInstance(). This hallucination occurs for lacking test-repair-oriented contexts (TROCtx) in the input causes. After providing the required contexts shown in Fig. 1, either including the definition of getDefaultInstance in class MountPOptions or the usage change of focal method in other production code, LLMs can generate the correct repaired test.
When developers are asked to repair test for this case, it is convenient for them to refer to the related contexts (TROCtx in Fig. 1) with the programming features (such as ‘goto definition’ and ‘find usages’) provided by Language Servers in IDEs. Inspired by the practices, Synbciatr constructs TROCtx by simulating developers’ behaviors in IDEs. Specifically, Synbciatr collects related contexts by directly interacting with Language Server, and filters out unnecessary ones based on Neural Rerankers. Finally, Synbciatr uses LLMs to repair obsolete test cases with constructed TROCtx.
III Methodology
In this section, we first introduce the overall framework of Synbciatr. Then we demonstrate all the types of TROCtx identified by Synbciatr. Finally, we will describe the main modules in Synbciatr in detail.
III-A Overview
The overall pipeline of Synbciatr is depicted in Fig. 2. Given the change of focal method and the obsolete test method as inputs, Synbciatr consists of three major steps. (1) Collecting TROCtx: Synbciatr analyzes all the related contexts from inputs and requests Language Server to collect and process them into candidate chunks; (2) Reranking TROCtx: Synbciatr reranks candidate chunks with queries constructed from the inputs; (3) Generating full prompt: Synbciatr aggregates the inputs and final TROCtx to generate the full prompt, which is used to repair test with LLM. We also provide a detailed overview of Synbciatr in Fig. 3.
III-B Types of TROCtx
Based on the characteristics of the test-repair task, we categorize TROCtx into three types: class contexts (ClassCtx), usage contexts (UsageCtx), and environment contexts (EnvCtx). In alignment with real-world developers’ practices, these categories of contexts can provide comprehensive and sufficient contextual information for test repair.
ClassCtx: the member accesses of a specific class type, including method accesses and field accesses. These contexts indicate the accurate operations supported by a given class type, serving to alleviate the hallucination of LLMs. Since new class types are introduced in parameter types and the return type, ClassCtx will be collected specially for ParamSynBC and RetSynBC. For example, the ClassCtx for the case caused by ParamSynBC in Fig. 1 is partly demonstrated in Listing ‣ III-B.
{listing}[!ht] {minted} [ frame=single, framesep=2mm, baselinestretch=1, fontsize=, breaklines=true ] java // methods are simplified as signatures … public static final int READONLY_FIELD_NUMBER = 1; public boolean hasReadOnly(); public static MountPOptions getDefaultInstance(); … ClassCtx of the new class MountPOptions.
UsageCtx: the changes of usages for the focal method. Usages of the updated focal method in other parts of the repository illustrate how to correctly call the method in the associated test. UsageCtx will be collected for all the types of SynBC. For example, the UsageCtx for case introduced in Fig. 1 is partly demonstrated in Listing ‣ III-B.
{listing}[!ht] {minted} [ frame=single, framesep=2mm, baselinestretch=1, fontsize=, breaklines=true ] diff … - mount(alluxioPath, ufsPath, MountOptions.defaults() + mount(alluxioPath, ufsPath, MountPOptions.getDefaultInstance() … UsageCtx of the focal method mount().
EnvCtx: the environmental changes of the focal method and the test method. Given a method , we define the class containing it and its parent classes as the environment of . These contexts indicate modifications of related identifiers and similar change patterns that exist externally to the method. EnvCtx will be collected for all types of SynBC. For example, the EnvCtx of the focal method for the case introduced in Fig. 1 is partly demonstrated in Listing ‣ III-B.
{listing}[!ht] {minted} [ frame=single, framesep=2mm, baselinestretch=1, fontsize=, breaklines=true ] diff … - return openFile(path, OpenFileOptions.defaults()); + return openFile(path, OpenFileOptions.getDefaultInstance()); … EnvCtx of of the focal method mount().
Synbciatr aims to collect and rerank different types of TROCtx respectively. In the following subsections, we will describe the details of how each type of TROCtx is collected, reranked, and ultimately aggregated as summarized in Fig. 3.
III-C Collecting TROCtx
To collect repo-level contexts, Synbciatr interacts with Language Servers via LSP [19]. Every request message conforms to the JSON-RPC-based protocol, which consists of the request type (such as ‘goto definition’) and the cursor position of the request identifier, which consists of the file path, and the indices of the line and column of the identifier within the file.
For every type of TROCtx, Synbciatr first automatically analyzes the inputs and locate key identifiers (identifiers that need to find definitions and references by LSP) with their positions using the static parser tree-sitter111https://tree-sitter.github.io/tree-sitter/. The required contexts are further collected through the Language Server using the python library multilspy222https://github.com/microsoft/monitors4codegen.
III-C1 Collecting ClassCtx
Based on the syntactic diff of the focal method, Synbciatr first identifies the new class types introduced in the method signature of the updated version (both parameter types and the return type will be checked). These new class types are considered as key identifiers for collecting ClassCtx. Through requesting Language Server, Synbciatr collects the definitions of these new class types as well as their parent classes. To split the collected contexts into identical chunks, Synbciatr cleans up all comments and the definition body is divided into member access operations. Only non-private declarations of fields and methods are collected. Specifically, for classes and private field accesses with Lombok annotations333https://projectlombok.org/features/(@Data, @GETTER, @SETTER), Synbciatr also retains related private field accesses.
III-C2 Collecting UsageCtx
For UsageCtx, the name identifier of the updated focal method is considered as key identifier. The collecting procedures are described in Algorithm 1. On line 3, Synbciatr locates the name identifier of the updated focal in the repository. On line 4, all the usages of the updated focal method are fetched by requesting Language Server. For each usage, Synbciatr collects the usage-diff texts (diffs that contain the change of usage) in two steps, generating the diff (lines 6- 10) and gathering the required contexts (lines 11- 18). Synbciatr formats the files that contain usages before generating diff to avoid missing information. Also, since usage-diff of invocation is not sufficient for ParamSynBC and RetSynBC, Synbciatr collects additional backward and forward contexts respectively (lines 13 and 16).
For the case in Fig. 1, Synbciatr finds ten usages in the repository. After filtering out usages in comments and repeated ones, four usage-diff texts are collected, one of which is demonstrated in Listing ‣ III-B.
III-C3 Collecting EnvCtx
EnvCtx is constructed with additional contexts indicating environmental changes, namely env-diff texts. Synbciatr collects EnvCtx for both focal method and test method. The class file to which the method belongs and its parent classes are treated as the method’s environment. Here, the class name identifiers are the key identifiers, which are used to find out the related parent classes and collect their diffs as env-diff texts. Also, diffs are generated after cleaning comments and reformatting.
III-D Reranking TROCtx
Synbciatr utilizes Neural Rerankers to filter and retain the most relevant contexts. As the test cases should be functionally consistent before and after the evolution, they have high similarity. Thus, our primary idea of rerankers is to make use of the original test to construct queries for TROCtx reranking. For the case in Fig. 1, we need to filter out unrelated member accesses and precisely retain the required method declaration “MountPOptions.getDefaultInstance()” from the collected ClassCtx of the new class MountPOptions. Given the text “MountOptions.defaults()” as the query, we can rerank the ClassCtx to get the most similar APIs with Neural Rerankers. As shown in Listing III-D, the reranking result reveals that the required API “MountPOptions.getDefaultInstance()” is ranked with the highest score under the given query.
[!ht] {minted} [ frame=single, framesep=2mm, baselinestretch=1, fontsize=, breaklines=true ] java // methods are simplified as signatures public static MountPOptions getDefaultInstance(); // top1 public MountPOptions getDefaultInstanceForType(); // top2 … Reranked ClassCtx of class MountPOptions.
III-D1 Constructing Queries
Based on the above ideas, the quality of the query texts decides the reranking results. However, as we have three types of TROCtxs but their formats (or granularity) are different, Synbciatr reranks each type of TROCtx with distinct queries instead of using a general one. In terms of granularity for TROCtx, ClassCtx is a fine-grained one where each context is a specific member access operation within the corresponding class (Listing ‣ III-B), while UsageCtx and EnvCtx are coarse-grained ones whose contexts directly refer to the diff texts of statements (Listings ‣ III-B and ‣ III-B). Overall, Synbciatr constructs two kinds of corresponding queries.
Constructing operation queries. We define operations as member accesses to a given class. Operation queries are used to rerank method and field declarations in ClassCtx for new class types. Based on our insight, the operations used in the original test can be reused as queries, namely operation queries, to rerank ClassCtx. For new class types in ParamSynBC and RetSynBC, Synbciatr constructs fine-grained queries by extracting operations that could potentially be accessed in the repaired test as shown in Algorithm 2. Specifically, Synbciatr constructs operation queries for new parameter and return class types respectively.
To construct operation queries for collected ClassCtx in ParamSynBC, starting from the invocation of the focal method, Synbciatr extracts operations specifically related to the obsolete parameters of the original focal method, which is deleted in the given SynBC. On line 8, Synbciatr traverses the def-use chains [22] of obsolete parameters to extract directly used operations, including methods accesses and field accesses. These operations are considered as operation queries. Besides, since the obsolete parameters may be refactored to be set by the new parameter, Synbciatr adds these operations in the form as setxxx on lines 4- 6.
To construct operation queries for collected ClassCtx in RetSynBC, starting from the invocation of the focal method, Synbciatr applies forward propagatability analysis by traversing the references of the return object of the focal method. During the analysis, related operations of the return object are extracted as the operation queries.
For the focal method in Fig. 1, only the parameter types changed. The finally extracted operation queries for the newly added parameter class type MountPOptions are as follows.
: {‘MountOptions.defaults()’, ‘setOptions()’}
Constructing statement queries. Synbciatr constructs coarse-grained queries for reranking diff texts of statements. In our settings, diff texts are collected after reformatting so that every line in the diff contains a separate statement. Since we only keep changed parts in diff texts, we should extract obsolete statements from the original test as the query, namely statement query, to rerank diff texts in UsageCtx and EnvCtx.
As recent studies have shown that LLM specializes in code summarizing and understanding tasks [23], [24], Synbciatr uses LLM to extract obsolete statements from the original test. Following existing works [25], we adopt LLMs with Few-Shot Learning and Chain-Of-Thought prompting to identify obsolete statements with the syntactic change of the focal method and original test method as inputs. LLM is first asked to summarize the syntactic differences of the focal method, and then find the obsolete test statements. Both the SynBC analysis (in natural language) and obsolete statements (stmts) are collected as coarse-grained statement queries. Specifically, to rerank the EnvCtx of the focal method, obsolete test statements are unrelated. Therefore, we use the intermediate SynBC analysis of the focal method as a coarse-grained query to search for env-diff contexts about the SynBC.
For the example in Fig. 1, the extracted statement queries are demonstrated as follows.
SynBC analysis: “The method mount() has been updated to accept an object of type ‘MountPOptions’ instead of ‘MountOptions’ as its third parameter.”
obsolete stmts: “MountOptions mountOptions = MountOptions.defaults(); mFileSystem.mount(alluxioPath, ufsPath, mountOptions;”
III-D2 Reranking with Queries
Synbciatr reranks each type of TROCtx based on different queries as shown in Fig. 3. For every reranking process, the top three most relevant contexts are retained by default.
To rerank ClassCtx, first, all the constructor declarations in ClassCtx are retained since they are necessary to construct instances for the class. Then, for other method and field declarations, Synbciatr transforms method declarations in ClassCtx into their signatures before reranking. According to the SynBC, ClassCtx of new parameter and return class type are reranked with and respectively.
To Rerank UsageCtx, every usage diff in UsageCtx is divided into original usage (by removing added lines in diffs) and updated usage (by removing deleted lines in diffs). Synbciatr reranks both the original and updated usages with obsolete statements as query, in which the maximum reranking score determines the relevance of the given usage diff.
To Rerank EnvCtx, EnvCtx of the focal method is directly reranked with SynBC analysis while EnvCtx of the test method is directly reranked with obsolete statements.
III-E Generating Full Prompt
The full prompt consists of the unified diff of the focal method, the original test, and the related contexts (TROCtx).
Following existing works using LLM for code tasks [26], Synbciatr cleans comments in codes to avoid influencing inferring program intentions. Then, Synbciatr generates the unified diff of the focal method after formatting.
The general structure of the final prompt is demonstrated in Fig. 3. Specifically, we start by setting up the system role of LLM as an expert in Java software evolution and briefly describe the test-repair task caused by syntactic changes of the focal method. Then, we introduce the original test that needs to be repaired followed by the unified diff of the focal method. At last, constructed TROCtx is listed into different types and we ask LLM to treat these contexts as references. For each group of contexts, we also provide the basic description. For the case in Fig 1 as an example, we add a description (‘Defined in class MountPOptions’) for ClassCtx of MountPOptions.
Finally, Synbciatr requests LLM with the generated full prompt to generate the repaired test.
IV Experimental Setup
Before the evaluation, in this section, we first introduce the construction of our benchmark dataset and then describe the baselines and metrics used in the evaluation. Finally, we will briefly show the implementation settings of our approach.
IV-A Benchmark Datasets
In this study, we focus on repairing obsolete test cases caused by syntactic changes in the focal method. To the best of our knowledge, Ceprot is the first and SOTA learning-based approach to co-evolve test cases at the method level [12]. Therefore, we reuse and refine the dataset provided by Ceprot, in which all the samples are collected from the top 1,500 Java projects on GitHub by Liu et al. [27].
The evaluation dataset, i.e., testing dataset, of Ceprot contains 520 samples extracted from 128 real-world Java projects. Based on that, we first filter out that samples with no syntactic changes in the focal method. After this step, 211 samples remain. Then, to ensure the high quality of the dataset, we manually filter noisy samples and augment new samples referring to the corresponding repository commit.
Filtering noisy samples. Samples with the following characteristics are filtered out as noises in the dataset:
the test method is incomplete or not yet implemented;
the focal method is incomplete with only the signature;
the focal method is not used in the given test.
We manually resolve the problems above by extracting correct co-evolution pairs with SynBC from the repository.
Augmenting new samples. To improve the diversity and generality of the dataset, we mine new samples to enrich the dataset. Specifically, we augment at most 2 samples from the real-world repository with different changing patterns for each commit in the original dataset.
To sum up, the final benchmark dataset consists of 137 samples from 54 projects. As shown in Fig. 4, the samples are diverse in the types of associated SynBC.
IV-B Baselines
To assess the effectiveness of Synbciatr, we consider two baselines from existing studies and our investigation, namely Ceprot and Naivellm.
Ceprot is the SOTA learning-based approach in updating obsolete test cases. According to Hu et al. [12], Ceprot is built on a code language model fine-tuned from CodeT5 and outperforms previous techniques. Taking notes that Ceprot needs edit sequences of the focal method as input, we reproduce it with clang-format444https://clang.llvm.org/docs/ClangFormat.html and difflib555https://docs.python.org/zh-cn/3/library/difflib.html. With the replication package of Ceprot, we retrained the model and saved the best checkpoint with the highest F1. According to the result of our replication (accuracy: 11.9%), the performance is consistent with the statistics in the paper (accuracy: 12.3%) [12].
Naivellm is developed based on LLM without related contexts. Compared with CodeT5, LLM is trained on huge sets of data with a much larger number of parameters. Therefore, LLM is more intelligent in general code tasks. In Naivellm, we directly use LLM to repair the test case with the focal diff (the unified diff of original and updated focal methods) and the original test method as input. Following the pre-processes in Synbciatr, comments in codes are cleaned and all the filtered codes are reformatted to a standardized style.
IV-C Metrics
Based on previous studies [12] [28], we design metrics to measure the quality of repaired test cases from two aspects, and finally get five metrics.
Textual Match. We use three specific metrics to measure how the generated code is close to the ground truth. (1) CodeBLEU. Developed from the classical machine translation evaluation metric BLEU [29], CodeBLEU [30] is widely used in evaluating code generation tasks by measuring the similarity with code semantics. Thus, we use CodeBLEU to assess the similarity between the generated method and the ground truth. (2) DiffBLEU. As only a part of statements are modified, we design the metric DiffBLEU to concentrate on the modified ones specifically. To compute it, we calculate the BLEU score of the repaired test whose unchanged statements are removed. DiffBLEU serves as a complement to CodeBLEU. (3) Accuracy (%). This metric represents the percentage of samples where the generated method is identical to the ground truth (exactly match).
Intent Match. Besides textual match metrics, we also design two other specific metrics to measure whether the generated repaired test cases are correct. Ideally, test repair should not change the intent of the test case, which means that the repaired test should keep the same assertions and input values. Referring to the original test and the ground truth respectively, we use two specific metrics to evaluate the repairability. (1) (%). represents the repairability referring to the original test. Specifically, we calculate as the percentage of samples where the generated test can be successfully compiled and shares the same assertions and input values with the original test. (2) (%). Similar to , is calculated referring to the ground truth. For each repaired test case, two developers with more than three years of Java programming experience are asked to perform manual evaluation separately. If they can’t reach an agreement, they will have a discussion on uncertain samples until they agree on consistent conclusions.
In addition, we also record the Syntax Pass Rate (SPR) and Compilation Pass Rate (CPR) as the success rate of syntax checking and compilation for generated codes respectively.
IV-D Implementation Settings
Synbciatr is built based on LangChain [31], which is a framework designed to simplify the procedures of developing applications powered by LLMs. With the APIs provided by LangChain, Synbciatr utilizes the SOTA LLM (GPT-4) [32] developed by openai and the widely-used open-source Neural Reranker (bge-reranker-v2-m3) [33] released by BAAI [34]. To reduce the impact of randomness, for each case in the dataset, we request LLM with the temperature as 0.1 three times and keep the best one for both Naivellm and SynBCIATR in evaluation.
V Evaluation
Based on the constructed benchmark dataset, we evaluate the effectiveness of our approach and conduct a comprehensive analysis of the results. We address these research questions:
RQ1:(Effectiveness of Synbciatr) Can Synbciatr effectively repair obsolete tests caused by SynBC?
RQ2:(Effectiveness of TROCtx) To what extent does the TROCtx contribute to correctly repairing the test?
RQ3:(Failure Analysis) Under which cases does Synbciatr fail to repair?
RQ4:(Efficiency) What is the efficiency of Synbciatr?
V-A RQ1: Effectiveness of Synbciatr
V-A1 Basic Evaluation on Effectiveness
To evaluate the performance of Synbciatr compared with baselines, we adopt different approaches to repair obsolete test cases in the benchmark dataset.
First, we conduct syntax validation on the generated test codes. As shown in Tab. I, the Syntax Pass Rate (SPR) of Ceprot is 47.8%, while LLM-based approaches all pass the validation (100%), which means that LLM demonstrates a higher proficiency in producing syntactically accurate code than CodeT5. Then we use the three textual metrics to measure the average similarity between the generated test code and the ground truth. The column Textual Match in Tab. I demonstrates that our approach outperforms baselines in terms of all three metrics. Specifically, LLM-based approaches have average improvements of 12.4% and 72.4% in terms of CodeBLEU and DiffBLEU compared with Ceprot, which indicates that LLM is more capable of understanding the semantics of code and generating repaired tests. With the highest scores on CodeBLEU and DiffBLEU, Synbciatr is also able to accurately repair 32.4% of the test cases, which also achieves varying degrees of improvement over CEPROT (4.4%) and NAIVELLM (28.7%).
| Approach | SPR(%) | Textual Match | ||
| CodeBLEU | DiffBLEU | Accuracy(%) | ||
| Ceprot | 47.8% | 73.5 | 26.3 | 4.4% |
| Naivellm | 100% | 81.9 | 44.0 | 28.7% |
| Synbciatr | 100% | 83.3 | 46.7 | 32.4% |
V-A2 Human Evaluation on Effectiveness
Since textual metrics focus on measuring the similarity of tokens and AST structure for the given codes, they can not well represent the correctness of repair. To bridge this gap, in this part, we replace the obsolete test cases with the generated repaired ones and manually check whether the replaced test is correct or not, where a correct repair should pass the compilation and keep the intent of the test unchanged.
| Approach | CPR(%) | Intent Match | |
| Ceprot | 33.3% | 19.9% (27) | 25.7% (35) |
| Naivellm | 88.5% | 69.1% (94) | 82.4% (112) |
| Synbciatr | 96.2% | 75.0% (102) | 90.4% (123) |
The column CPR in Tab. II represents the success rate of compilation after repairing the original test case in the repository. As a result, Synbciatr fixes the most compilation errors caused by SynBC. In terms of the two metrics of repairability, Synbciatr both outperforms the baselines. As shown in the column Intent Match, we can observe that Synbciatr correctly repair 75.0% (102/136) and 90.4% (123/136) test cases in alignment with the intent of the ground truth and the original test respectively.
As for specifically, Synbciatr achieves improvements of 248.6% and 9.8% when compared to Ceprot and naiveLLM respectively.
V-B RQ2: Effectiveness of TROCtx
In this research question, we focus on comparing our proposed Synbciatr with Naivellm to evaluate the effectiveness of TROCtx. As introduced in Section I, TROCtx serves as contextual information to reduce hallucinations during the generation of LLM (LLM below all refers to GPT-4).
According to existing work [35] and our investigation, we define hallucinations as using undefined methods, variables, or classes in the LLM-generated codes. Based on the results of human evaluation, we manually identify hallucinations and summarize them into two types: common hallucinations and outdated hallucinations.
Common hallucinations. Common hallucinations are caused by the lack of direct contextual information from the change of focal method signature. For the case in Fig. 1, a new class (MountPOptions) is set as the third parameter type. Therefore, LLM falls into common hallucination without the required class contexts of MountPOptions.
Outdated hallucinations. Outdated hallucinations are caused by the lack of implicit contextual information from the focal method body. For the example in Fig. 5, the method findUnknown used in the test is refactored to be accessed from an instance of class BitstreamFormat. Without this specific knowledge, the test generated by LLM still uses the method in an outdated way, which causes hallucinations.
To assess the effectiveness of TROCtx constructed by Synbciatr, we collect all the hallucinations for Naivellm and Synbciatr. As shown in Tab. III, hallucinations occur in both approaches. For the 6 common hallucinations that Naivellm fails, Synbciatr fixes all of them and correctly generates repaired tests with related contexts, indicating that Synbciatr has the general capability to precisely and effectively collect required contexts from the syntactic change of the focal method. Besides, Synbciatr also reduces outdated hallucinations by 25%, in which two cases are fixed with the constructed TROCtx (UsageCtx and EnvCtx specifically).
| Approach | Common Hall. | Outdated Hall. | Total |
| Naivellm | 6 | 8 | 14 |
| Synbciatr | 0 (-100.0%) | 6 (-25.0%) | 6 (-57.1%) |
V-C RQ3: Failure Analysis
Synbciatr is designed to repair the original test case with its intent unchanged. Considering that the ground truth may contain semantic changes such as adding new assertions and altering input values, we further investigate the 13 cases that fail in terms of according to Tab. II. Finally, we summarize three reasons for the failure.
Uses Unimported Classes. For four cases, LLM directly uses classes that are not imported in the file of the test method when generating codes. The key reason is that we focus on collecting contexts by analyzing the syntactic change of the focal method at the method level. Without analyzing the focal class and changes in the methods, the constructed contexts by Synbciatr may be inadequate and result in failure. Fortunately, this type of failure is easy to fix.
Complex Focal Changes. For eight cases, LLM fails to correctly generate the expected test with limited contexts as the changes of focal methods in class-level or repo-level upgrades are complex. Specifically, the complex focal change leads to outdated hallucinations and incorrect invocations. This type is challenging to resolve because of the difficulty of collecting implicit contexts and the limited capability of current LLMs.
Fails to construct TROCtx. Specifically for one case, We find that Language Server fails to provide intelligent features for repositories containing configuration error, which results in the failure of Synbciatr to construct TROCtx. Overall, it is hard even for developers to manually repair tests in these repositories, in which they can not infer related contexts in IDEs either.
V-D RQ4: Efficiency
Compared to Naivellm, Synbciatr adds the steps to construct TROCtx. When querying LLM to repair the test, the prompt of Synbciatr is longer with the constructed TROCtx. In this research question, we evaluate the time efficiency and the token count of Synbciatr compared to Naivellm.
In particular, we divide the process of Synbciatr into constructing TROCtx and querying LLM for repair. As shown in Fig. 6, the time cost of Synbciatr mainly depends on constructing TROCtx, in which the response time of LLM API of two approaches is close. We notice that the Language server may fail to synchronize old repositories due to connection timeout for missing dependencies, which results in the outliers with much longer time. In our design, Synbciatr trades time for accuracy. Besides, the average time of Synbciatr falls into 10s to 30s, which is acceptable for developers since test repair is not an interactive task.
In terms of the token cost of Synbciatr, we observe that the tokens of the prompt roughly doubled in number after being augmented with TROCtx from Fig. 7. This evidence also indicates that the TROCtx constructed by Synbciatr is precise.
VI Threats to Validity
In this section, we discuss the possible threats to validity from external and internal aspects respectively.
External Validity. The main threats to external validity come from the evaluation dataset. We reuse the existing dataset, which may not be representative of all possible real-world syntactic changes, so we check and augment the dataset to collect diverse samples from related commits. Besides, it is inevitable to avoid data leakage from popular open-source projects in GitHub. These projects widely use standard design patterns [36], which can be learned by LLM during training. Although experimental results show that Naivellm correctly repairs some of the tests with data leakage, it does not affect the effectiveness evaluation of Synbciatr. According to the analysis in RQ2, Synbciatr outperforms Naivellm in cases where Naivellm fails for hallucinations.
Internal Validity. In our experiments, a major threat to internal validity is the possible bias in human evaluation. To mitigate it, we invite two senior developers to manually verify the generated tests and annotate explanations. The final result is collected after they reach an agreement after discussion.
VII Related Work
VII-A Production-test Co-evolution
Production-test co-evolution refers to co-evolving the test codes with the changes in production codes. Most of the previous studies focus on identifying production-test co-evolution pairs [37], [6], [7], [11]. Recently, more works adopt learning-based techniques to automatically identify obsolete tests [10], [11]. While these works focus on identification only, our approach targets automatically repairing the obsolete test cases directly to relieve the burden of developers.
There are two types of obsolete test cases that need repair, one is the GUI-oriented event sequence test case, the other is the code-oriented method test case. Nowadays, many researchers studied the automated repair of GUI test cases, especially on Android applications [38], [11]. However, for code-oriented method test cases, there are limited studies focused on repairing code-oriented method test cases automatically. The biggest difference between them is that the searching space for the GUI-oriented test event is more oblivious, which can be obtained by traversing the related GUI widget trees, while the code-oriented method test case has a larger searching space in the whole repository.
For code-oriented method test case repairing, Hu et al. [12] proposed the first transformer-based approach to update obsolete tests with two stages, identifying and updating, which is the SOTA work in this area. However, it is based on pre-trained models with fewer parameters and lacks contextual information. Compared to this, our work uses larger language models (LLMs) with automatically constructed contexts from the repository.
VII-B LLM-based Code Generation
Automated code generation with LLM brings huge improvements in production efficiency. Repo-level code generation represents the task of generating codes based on a broader context of the repository [39]. It is challenging due to the lack of domain-specific knowledge, which results in hallucination [40], [41]. Several studies leverage RAG to improve the performance of LLM in specific code tasks by providing similar codes or results into the query prompt [39], [25]. However, the contexts are retrieved based on simple metrics such as textual similarity without code semantics. Recently, some works have focused on improving the capability of LLM by static analysis [42], [18]. Specifically, monitor-guided decoding (MGD) [18] is a novel approach to bring IDE-assistance from developers to LLMs to guide the decoding when generating codes, in which IDE-assistance is the intelligent features provided by Language Server. Compared to this work, we combine static analysis with RAG to provide more contexts instead of guiding the decoding process.
VIII Conclusions
We propose Synbciatr, an LLM-powered approach to automatically repair obsolete test cases caused by syntactic changes of the focal methods. The key idea of Synbciatr is to combine the static analysis and neural rerankers to precisely construct test-repair-oriented contexts from the updated repository. Experimental results show that Synbciatr’s effectiveness outperforms baseline approaches on both the textual- and intent-matching metrics. Besides, with the augmentation of TROCtx constructed by Synbciatr, hallucinations are reduced by 57.1%. The overall results demonstrate that adopting static analysis techniques to improve the capability of LLM yields excellent performance and could be extended to other code-related tasks.